# Toward a test of Gaussianity of a gravitational wave background

### v1

@@ II.4 One-point statistics
Figure 1 shows the ensemble distributions (5000 samples/simulations) of the sample statistics, mean (${\cal E}_{1}$), variance (${\cal V}_{2}$), skewness (${\cal S}_{3}$), and kurtosis (${\cal K}_{4}$), of the one-point function of pulsar timing residuals in 30 yr-PTA simulations with 100 pulsars, scattered anisotropically Babak et al. (2024).
@@
Figure 1: Ensemble distributions of the first four moments/one-point metrics of PTA timing residuals, given only GWB and red noise (RN) components; $n_{\rm psrs}=100$ pulsars and $n_{\rm sims}=5000$ simulations.
@@
Starting with the GWB signal, our first comment goes to the variance across frequency bins. This follows ${\cal V}_{2}\sim f^{-13/3}$ as shown, reflective of the spectrum corresponding to the input GWB. In relation, the fourth centralized moment, which is proportional to the kurtosis, admits ${\cal K}_{4}\sim\left(Af^{-13/3}\right)^{2}$. This is owed to the Gaussian nature of the GWB, such that the higher moments of the underlying field factorize into a product-sum of two-point functions. In the astrophysical setting, with an ensemble of SMBHBs as the source of GWB, this relation in the one-point function can be expected to manifest in the cumulants of the spectrum, as discussed in Lamb and Taylor (2024). We also highlight that while the distribution of the mean is fairly consistent with a Gaussian distribution, the distribution of the skewness unequivocally deviates from this trend, consistent with our results based on simulating GWB maps and computing their one-point statistics. Our results show that the higher moments of the PTA residuals tend to a non-Gaussian ensemble distribution, attributed to a large scale quadrupolar correlation characteristic of a GWB. In particular, we find that the presence of large scale correlation implies that the central limit theorem cannot be applied. See Appendix A.
@@
However, we find that the dominant red noise component in the pulsars similarly produce a non-Gaussian ensemble distribution of sample statistics, although for a completely different reason compared with the GWB component. The variation of the red noise spectrum across pulsars is sourcing a huge departure of the resulting ensemble distribution away from a Gaussian distribution. This is notably more pronounced by orders of magnitude compared with the GWB, particularly for the higher moments of the pulsar timing residuals’ one-point function. It is worth highlighting that the dominant red noise also completely contaminates the one-point statistics, as can be seen in the magnitudes of the variance and the kurtosis. When both contributions are considered in the simulations, among other usual components such as white noise Roebber (2019); Vallisneri et al. (2020), the resulting ensemble distributions do not turn away too far off from that of the red noise component alone. Adding more pulsars to the observation can be expected to only amplify the induced non-Gaussianity due to the variation of the red noise spectrum. Although one might expect the sum of uncorrelated random variables to approach a Gaussian distribution asymptotically, the noise is not identically distributed and different pulsars can exhibit orders of magnitude differences in their noise contributions. The cumulants of the red noise is dominated by a small number of pulsars with large noise components.
@@
This illustrates that both the GWB and red noise components source ensemble distributions that stray far from that of a Gaussian distribution, for the distinct reasons discussed above. For the skewness, the ensemble distribution in both GWB and RN cases turns out to be narrower compared with a corresponding Gaussian one, but with a long tail. However, it must be emphasized that the sample statistic for red noise also varies by orders of magnitude compared with that of GWB. The same holds for the kurtosis, which manifests orders of magnitude difference in the red noise compared with GWB. This can be attributed to the variation in the red noise spectrum across pulsars, and can be tested by varying the number of pulsars. Note that the plots are shown in log-scale to make the differences perceivable by eye. To support of the above statements that were drawn from visual inspection, we also performed standard normality tests d’Agostino (1971); d’Agostino and Pearson (1973); Shapiro and Wilk (1965); Phipson and Smyth (2010); Panagiotakos (2008); Virtanen et al. (2020) that assign $p$-values relative to a null hypothesis that a sample comes from a normal/Gaussian distribution. For the variance, skewness, and kurtosis, in all frequency bins, the results have unambiguously ruled out the null hypothesis ($p\ll 0.05$), implying that the samples presented in Figure 1, hence also Figure 2, do not conform to a Gaussian ensemble distribution.
@@
III Two-point statistics
In this section, we lay down our results on two-point function statistics of the GWB signal. We also consider the case when both signal and noise are added to the observation.
The mean statistic, ${\cal E}_{1}$, of the timing residual cross correlation of a pulsar pair $a$ and $b$ due to an isotropic and Gaussian GWB (5) can be shown to be
${\cal E}_{1}[r_{a}r_{b}]=\tilde{a}^{2}c_{ab}\,,$ (7)
@@ III Two-point statistics
${\cal V}_{2}[r_{a}r_{b}]=\tilde{a}^{4}\left(1+c^{2}_{ab}\right)\,,$ (8)
${\cal S}_{3}[r_{a}r_{b}]=2\tilde{a}^{6}c_{ab}\left(3+c_{ab}^{2}\right)\,,$ (9)
@@
${\cal K}_{4}[r_{a}r_{b}]=3\tilde{a}^{8}\left(3+14c_{ab}^{2}+3c_{ab}^{4}\right)\,,$ (10)
@@
respectively. Note that all of the above higher moments of the two-point function reduce to products of $c_{ab}$, by virtue of the Gaussianity of the underlying field. For this case, Gaussianity implies that the higher moments of the timing residual correlation give no more further information beyond the HD curve Akrami et al. (2020). Since a GWB signal is differentiated from noise by spatial correlation, Gaussianity can be tested directly by measuring the higher moments of the two-point function, since only the signal can generate coherent correlations.
@@
Figure 3: The signal of an isotropic, Gaussian nanohertz GWB in noise-free 30 yr-PTA simulations in the first frequency bin (100 pulsars, and 300 realizations). The theoretical curves are given by (7-10). The results in the other frequency bins are similar, with relevant amplitudes scaled accordingly, $\tilde{a}^{2}\rightarrow\tilde{a}^{2}(f/f_{1})^{-13/3}$, where $\tilde{a}^{2}$ and $f_{1}$ are the correlation amplitudes and the frequency in the first bin, respectively.
@@
To support the discussion, we simulate 30 yr-PTA realizations with one hundred pulsars, injected with a HD-correlated GWB with an amplitude $A_{\rm gw}=2.4\times 10^{-15}$ and circular SMBHB spectral index $\gamma_{\rm gw}=13/3$ Phinney (2001). We then pair up the pulsars in angular separation bins for every pair in a PTA realization, and then in each angular bin obtain the sample statistic (mean, variance, skewness, and kurtosis) of $\alpha_{ak}\alpha_{bk}$ and $\beta_{ak}\beta_{bk}$ over the pulsar pairs ($a,b$) and frequency bins, $f_{k}=k/T$. This is repeated $300$ times to create an ensemble, displaying the mean and the cosmic variance of the two-point function’s sample statistics.
@@
The theoretical description (7-10) is confirmed by our noise-free simulations, as in Figure 3, obtained by injecting only the GWB signal into the pulsars. The results also confirm that the signal in the higher frequency bins are described simply by the same signal in the first bin, with the amplitudes rescaled accordingly to the power spectrum of the signal, $\tilde{a}^{2}\rightarrow\tilde{a}^{2}(f/f_{1})^{-13/3}$, where $\tilde{a}^{2}$ and $f_{1}$ are the correlation amplitudes and the frequency in the first bin, respectively. This shows that the higher moments of a Gaussian GWB’s two-point function in a PTA can be described by the spectrum and spatial correlation that is already associated with its mean.
@@
Figure 4: The timing residual correlation two-point statistics in 30 yr-PTA simulations in the first frequency bin (100 pulsars, and 300 realizations) with varying degrees of noise in the pulsars. Blue is noise-free (GWB signal); green is with soft noise (RN as loud as signal); red is with mild noise (RN one order of magnitude above signal). The theoretical curves (blue) are given by (7-10).
@@
Figure 5: The timing residual correlation two-point statistics in 30 yr-PTA simulations (same as in Figure 4) in the 14th frequency bin (100 pulsars, 300 realizations) with varying degrees of noise in the pulsars. Blue is noise-free (GWB signal); green is with soft noise (RN as loud as signal); red is with mild noise (RN one order of magnitude above signal). The theoretical curves (blue) are given by (7-10), with the amplitude rescaled as $\tilde{a}^{2}k^{-13/3}$.
@@
In the astrophysical PTA setting the signal is dominated by red noise in pulsars that has similar time scales as GWB timing residuals. This renders the one-point statistics unusable for testing the Gaussianity of the signal. To understand how this manifests in the two-point function, we can take an observation from a pulsar $a$ to be $r_{a}=s_{a}+n_{a}$ where $s_{a}$ is the GWB signal (satisfying (7-10)) and $n_{a}$ is the noise. In a pure-noise case, the analogous two-point function statistics can be obtained simply by replacing the HD correlation, $c_{ab}$, in (7-10) by $\delta_{ab}$. Our pure-noise simulations, opposite to the noise-free case, support this simple model of an uncorrelated noise in the higher moments of the two-point function. However, in the realistic scenario where both signal and noise are present in the observation, a more precise way to model an (effective) correlation is by expanding the powers of $r_{a}r_{b}\equiv(s_{a}+n_{a})(s_{b}+n_{b})$, subject to simplifications that apply such as Gaussian factorization. Continuing it this way is tedious, and can perhaps pay off as a way to include the information in the two-point function for GWB detection purposes. We shall leave this for future work. Nonetheless, our simulations can tease the features to expect in the two-point function when noise is present as well as the signal, as shown in Figures 4 and 5 in the 1st ($f_{1}\sim 1.06$ nHz) and 14th ($f_{14}\sim 14.8$ nHz) frequency bins of the data.
@@
In the analysis, we consider several levels of the noise relative to the amplitude of the GWB signal ($A_{\rm gw}=2.4\times 10^{-15}$). First is the noise-free level, for which case the ensemble uncertainty can be identified with the cosmic variance Roebber et al. (2016); Roebber and Holder (2017); Allen (2023); Bernardo and Ng (2022). Next one is ‘soft’ noise, when the red noise spectrum in the pulsars in each realization is drawn with amplitudes $\log_{10}A_{\rm rn}={\cal U}(-17,-15)$ and spectral shape indices $\gamma_{\rm rn}={\cal U}(2,6)$, resulting to the overall noise becoming only as loud as the signal in the observation. We also consider a ‘mild’ noise case when instead the noise can be drawn up to an order of magnitude above the signal, $\log_{10}A_{\rm rn}={\cal U}(-17,-14)$, and a ‘loud’ noise scenario $\log_{10}A_{\rm rn}={\cal U}(-17,-13)$ that meets the conditions we considered when discussing one-point statistics. Figures 4-5 illustrate the resulting ensemble statistics of the higher moments of the timing residual correlation for the noise-free, soft-, and mild-noise scenarios. Note that the same signal in Figure 3 is shown in Figures 4-5 (blue points traced by the theoretical (7-10)), except with the mean and the skewness expressed in log-scale for visual clarity when the noise is taken into account.
@@
The results show that despite being spatially-uncorrelated and drawn from independent Gaussian random processes the red noise remains to dominate the moments of the two-point function in the first frequency bin, particularly, in the variance and the kurtosis, even in the soft and mild PTA noise scenarios we considered. This domination of the noise is only amplified in the loud noise case. Nonetheless, the figures suggest that even with only one hundred pulsars (consistent with present PTA data) the mean of the two-point function can already hint toward a detection of the signal, which may be followed up by searching for the signal of Gaussianity (7-10) in the higher moments. This is reminiscent of the present state-of-the-art PTA science, indicating a compelling evidence of the sought GWB signal.
@@
We have confirmed the same conclusions on two-point function statistics with 200 and 400 pulsars. In this work we foocus on simulations with 100 pulsars because that is the current state-of-the-art in the field.
IV Outlook
@@
Our analysis has no assumption about the source statistics, but it will be important to draw this connection at some point Allen and Valtolina (2024). It remains to apply our tests to present data and forecast results expected in the forthcoming PTA/SKA precision era Lazio (2013); Weltman et al. (2020); Chandra Joshi et al. (2022); Çalışkan et al. (2024), which we will defer to future work. A Gaussian signal may turn out to be more significant in real data than is implied here, even though we suspect that the non-Gaussianity due to the noise will continue to dominate the cumulants. This is because the intrinsic noise parameters of each pulsar are measured, to an extent, post-sampling in the standard PTA search pipeline for a common spatially-correlated signal. The constrained noise parameters, particularly of the louder pulsars, may eventually be used to reduce their influence on the cumulants in both one- and two-point statistics. Real data is of course trickier, and there are plenty of challenges with PTA data that must be dealt with such as the practical limits to the sensitivity curve Hazboun et al. (2019), and the fact that the timing data contains gaps.
On the other hand, the statistical properties of a Gaussian GWB in the one- and two-point function discussed in this work are general and independent of any particular primordial model. The non-Gaussianity in the one-point cumulants (left of Figure 2) and the two-point signal (7-10) are tied to the quadrupolar spatial correlation, due to the tensorial nature/polarization of GWs. This implies that analogous tests may be setup to study GWB in other GW bands, such as for future space based GW detectors that are expecting to meet a foreground GWB from white dwarfs, among other relevant cosmological GWBs motivated in theory. The formalisms drawn out in this work are easily extendible to accommodate non-Einsteinian and subluminal GW propagations Chamberlin and Siemens (2012); Qin et al. (2021); Arzoumanian et al. (2021b); Chen et al. (2021, 2022); Wu et al. (2023); Bernardo and Ng (2023b, c, d, 2024); Liang et al. (2023); Cordes et al. (2024). When the tests come to fruition, it will be inevitably important to relate constraints on non-Gaussianity in the GWB observables to theory, requiring predictions beyond the power spectrum Powell and Tasinato (2020); Tasinato (2022); Zhu (2023). We leave this to future work.
Acknowledgements.

### v2

@@ 2.4 One-point statistics
Figure 1 shows the ensemble distributions (5000 samples/simulations) of the sample statistics, mean (${\cal E}_{1}$), variance (${\cal V}_{2}$), skewness (${\cal S}_{3}$), and kurtosis (${\cal K}_{4}$), of the one-point function of pulsar timing residuals in 30 yr-PTA simulations with 100 pulsars, scattered anisotropically [59].
@@
Figure 1: Ensemble distributions of the first four moments/one-point metrics of PTA timing residuals, given only GWB and red noise (RN) components; $n_{\rm psrs}=100$ pulsars and $n_{\rm sims}=5000$ simulations; red points/violins have both GWB and RN.
@@
Starting with the GWB signal, our first comment goes to the variance across frequency bins. This follows ${\cal V}_{2}\sim f^{-13/3}$ as shown, reflective of the spectrum corresponding to the input GWB. In relation, the fourth centralized moment, which is proportional to the kurtosis, admits ${\cal K}_{4}\sim\left(Af^{-13/3}\right)^{2}$. This is owed to the Gaussian nature of the GWB, such that the higher moments of the underlying field factorize into a product-sum of two-point functions. In the astrophysical setting, with an ensemble of SMBHBs as the source of GWB, this relation in the one-point function can be expected to manifest in the cumulants of the spectrum, as discussed in [75]. We also highlight that while the distribution of the mean is fairly consistent with a Gaussian distribution, the distribution of the skewness unequivocally deviates from this trend, consistent with our results based on simulating GWB maps and computing their one-point statistics. Our results show that the higher moments of the PTA residuals tend to a non-Gaussian ensemble distribution, attributed to a large scale quadrupolar correlation characteristic of a GWB. In particular, we find that the presence of large scale correlation implies that the central limit theorem cannot be applied. See Appendix A.
@@
However, we find that the dominant red noise component in the pulsars similarly produce a non-Gaussian ensemble distribution of sample statistics, although for a completely different reason compared with the GWB component. The variation of the red noise spectrum across pulsars is sourcing a huge departure of the resulting ensemble distribution away from a Gaussian distribution. This is notably more pronounced by orders of magnitude compared with the GWB, particularly for the higher moments of the pulsar timing residuals’ one-point function. It is worth highlighting that the dominant red noise also completely contaminates the one-point statistics, as can be seen in the magnitudes of the variance and the kurtosis. When both contributions are considered in the simulations, among other usual components such as white noise [76, 77], the resulting ensemble distributions do not turn away too far off from that of the red noise component alone. Adding more pulsars to the observation can be expected to only amplify the induced non-Gaussianity due to the variation of the red noise spectrum. Although one might expect the sum of uncorrelated random variables to approach a Gaussian distribution asymptotically, the noise is not identically distributed and different pulsars can exhibit orders of magnitude differences in their noise contributions. The cumulants of the red noise is dominated by a small number of pulsars with large noise components.
@@
This illustrates that both the GWB and red noise components produce a non-Gaussian distribution of the sample statistics across realisations ; for GWB, this is due to the large scale correlation, whereas for the red noise, this is due to the variation of the spectrum across pulsars. For the skewness, the ensemble distribution in both GWB and RN cases turns out to be narrower compared with a corresponding Gaussian one, but with a long tail. However, it must be emphasized that the sample statistic for red noise also varies by orders of magnitude compared with that of GWB. The same holds for the kurtosis, which manifests orders of magnitude difference in the red noise compared with GWB. This can be attributed to the variation in the red noise spectrum across pulsars, and can be tested by varying the number of pulsars. Note that the plots are shown in log-scale to make the differences perceivable by eye. To support of the above statements that were drawn from visual inspection, we also performed standard normality tests [78, 79, 80, 81, 82, 83] that assign $p$-values relative to a null hypothesis that a sample comes from a normal/Gaussian distribution. For the variance, skewness, and kurtosis, in all frequency bins, the results have unambiguously ruled out the null hypothesis ($p\ll 0.05$), implying that the samples presented in Figure 1, hence also Figure 2, do not conform to a Gaussian ensemble distribution.
@@
3 Two-point statistics
In this section, we lay down our results on two-point function statistics of the GWB signal. We also consider the case when both signal and noise are added to the observation. We add that the cumulants of the two-point function are operationally defined the same way as in the one-point case (third paragraph of Section 2.2), except that the input are now a pair of residuals/Fourier bins, e.g., $\alpha_{ak}\beta_{bk}$ with $a\neq b$, and sample averages/statistics are gathered over pulsar pairs in a realisation.
The mean statistic, ${\cal E}_{1}$, of the timing residual cross correlation of a pulsar pair $a$ and $b$ due to an isotropic and Gaussian GWB (2.5) can be shown to be
${\cal E}_{1}[r_{a}r_{b}]=\tilde{a}^{2}c_{ab}\,,$ (3.1)
@@ 3 Two-point statistics
${\cal V}_{2}[r_{a}r_{b}]=\tilde{a}^{4}\left(1+c^{2}_{ab}\right)\,,$ (3.2)
${\cal S}_{3}[r_{a}r_{b}]=2\tilde{a}^{6}c_{ab}\left(3+c_{ab}^{2}\right)\,,$ (3.3)
@@
${\cal K}_{4}[r_{a}r_{b}]=3\tilde{a}^{8}\left(3+14c_{ab}^{2}+3c_{ab}^{4}\right)\,,$ (3.4)
@@
respectively. Note that all of the above higher moments of the two-point function reduce to products of $c_{ab}$, by virtue of the Gaussianity of the underlying field. For this case, Gaussianity implies that the higher moments of the timing residual correlation give no more further information beyond the HD curve [60]. Since a GWB signal is differentiated from noise by spatial correlation, Gaussianity can be tested directly by measuring the higher moments of the two-point function, since only the signal can generate coherent correlations.
@@
Figure 3: The signal of an isotropic, Gaussian nanohertz GWB in noise-free 30 yr-PTA simulations in the first frequency bin (100 pulsars, and 300 realizations). The theoretical curves are given by (3.1-3.4). The results in the other frequency bins are similar, with relevant amplitudes scaled accordingly, $\tilde{a}^{2}\rightarrow\tilde{a}^{2}(f/f_{1})^{-13/3}$, where $\tilde{a}^{2}$ and $f_{1}$ are the correlation amplitudes and the frequency in the first bin, respectively.
@@
To support the discussion, we simulate 30 yr-PTA realizations with one hundred pulsars, injected with a HD-correlated GWB with an amplitude $A_{\rm gw}=2.4\times 10^{-15}$ and circular SMBHB spectral index $\gamma_{\rm gw}=13/3$ [26]. We pair up the pulsars in angular separation bins for every pair in a PTA realization, and then in each angular bin obtain the sample statistic (mean, variance, skewness, and kurtosis) of $\alpha_{ak}\alpha_{bk}$ and $\beta_{ak}\beta_{bk}$ over the pulsar pairs ($a,b$) and frequency bins, $f_{k}=k/T$. This is repeated $300$ times to create an ensemble, displaying the mean and the cosmic variance of the two-point function’s sample statistics. We add that in our analysis, all frequency and angular bins are treated independently. Optimal averages as in [84, 85] for two-point statistics will certainly be worth looking into for future work.
@@
The theoretical description (3.1-3.4) is confirmed by our noise-free simulations, as in Figure 3, obtained by injecting only the GWB signal into the pulsars. The results also confirm that the signal in the higher frequency bins are described simply by the same signal in the first bin, with the amplitudes rescaled accordingly to the power spectrum of the signal, $\tilde{a}^{2}\rightarrow\tilde{a}^{2}(f/f_{1})^{-13/3}$, where $\tilde{a}^{2}$ and $f_{1}$ are the correlation amplitudes and the frequency in the first bin, respectively. This shows that the higher moments of a Gaussian GWB’s two-point function in a PTA can be described by the spectrum and spatial correlation that is already associated with its mean.
@@
Figure 4: The timing residual correlation two-point statistics in 30 yr-PTA simulations in the first frequency bin (100 pulsars, and 300 realizations) with varying degrees of noise in the pulsars. Blue is noise-free (GWB signal); green is with soft noise (RN as loud as signal); red is with mild noise (RN one order of magnitude above signal). The theoretical curves (blue) are given by (3.1-3.4).
@@
Figure 5: The timing residual correlation two-point statistics in 30 yr-PTA simulations (same as in Figure 4) in the 14th frequency bin (100 pulsars, 300 realizations) with varying degrees of noise in the pulsars. Blue is noise-free (GWB signal); green is with soft noise (RN as loud as signal); red is with mild noise (RN one order of magnitude above signal). The theoretical curves (blue) are given by (3.1-3.4), with the amplitude rescaled as $\tilde{a}^{2}k^{-13/3}$.
@@
In the astrophysical PTA setting the signal is dominated by red noise in pulsars that has similar time scales as GWB timing residuals. This renders the one-point statistics unusable for testing the Gaussianity of the signal (Section 2.4). To understand how this manifests in the two-point function, we can take an observation/timing residual from a pulsar $a$ to be $r_{a}=s_{a}+n_{a}$ where $s_{a}$ is the GWB signal (satisfying (3.1-3.4)) and $n_{a}$ is the noise. In a pure-noise case, the analogous two-point function statistics can be obtained simply by replacing the HD correlation, $c_{ab}$, in (3.1-3.4) by $\delta_{ab}$. Our pure-noise simulations, opposite to the noise-free case, support this simple model of an uncorrelated noise in the higher moments of the two-point function. However, in the realistic scenario where both signal and noise are present in the observation, a more precise way to model an (effective) correlation is by expanding the powers of $r_{a}r_{b}\equiv(s_{a}+n_{a})(s_{b}+n_{b})$, subject to simplifications that apply such as Gaussian factorization. Continuing it this way is tedious, and can perhaps pay off as a way to include the information in the two-point function for GWB detection purposes. We shall leave this for future work. Nonetheless, our simulations can tease the features to expect in the two-point function when noise is present as well as the signal, as shown in Figures 4 and 5 in the 1st ($f_{1}\sim 1.06$ nHz) and 14th ($f_{14}\sim 14.8$ nHz) frequency bins of the data.
@@
In the analysis, we consider several levels of the noise relative to the amplitude of the GWB signal ($A_{\rm gw}=2.4\times 10^{-15}$). First is the noise-free level, for which case the ensemble uncertainty can be identified with the cosmic variance [68, 69, 70, 71]. Next one is ‘soft’ noise, when the red noise spectrum in the pulsars in each realization is drawn with amplitudes $\log_{10}A_{\rm rn}={\cal U}(-17,-15)$ and spectral shape indices $\gamma_{\rm rn}={\cal U}(2,6)$, resulting to the overall noise becoming only as loud as the signal in the observation. We also consider a ‘mild’ noise case when instead the noise can be drawn up to an order of magnitude above the signal, $\log_{10}A_{\rm rn}={\cal U}(-17,-14)$. Figures 4-5 illustrate the resulting ensemble statistics of the higher moments of the timing residual correlation for the noise-free, soft-, and mild-noise scenarios. Note that the same signal in Figure 3 is shown in Figures 4-5 (blue points traced by the theoretical (3.1-3.4)), except with the mean and the skewness expressed in log-scale for visual clarity when the noise is taken into account.
@@
The results show that despite being spatially-uncorrelated and drawn from independent Gaussian random processes the red noise remains to dominate the moments of the two-point function in the first frequency bin, particularly, in the variance and the kurtosis, even in the soft and mild PTA noise scenarios we considered. This domination of the noise is only amplified in the loud noise case. Nonetheless, the figures suggest that even with only one hundred pulsars (consistent with present PTA data) the mean of the two-point function can already hint toward a detection of the signal, which may be followed up by searching for the signal of Gaussianity (3.1-3.4) in the higher moments. This is reminiscent of the present state-of-the-art PTA science, indicating a compelling evidence of the sought GWB signal.
@@
We have confirmed the same conclusions on two-point function statistics with 200 and 400 pulsars. For these cases, the significance of the signal with respect to the noise in the two-point statistics has increased as expected, e.g., compared to 100 pulsars, in our mild noise scenario, the variance of the noise compared to the signal in the 14th frequency bin has decreased by 4% and 10% for 200 and 400 pulsars, respectively. Accurately modelling the noise (as described in the 5th paragraph of this section) should further enable a more precise determination of the signal’s properties via two-point statistics. In this work we focus on simulations with 100 pulsars because that is the current state-of-the-art in the field. The conclusions on one- and two-point statistics are also robust to the consideration of pulsar white noise, which is uncorrelated spatially and in time, and has negligible contribution. We leave for future work the analysis with dispersion measure/chromatic noises which depend on frequency of radio observation and may be relevant for some pulsars in a PTA.
4 Outlook
@@
Our analysis has no assumption about the source statistics, but it will be important to draw this connection at some point [70, 86]. It remains to apply our tests to present data and forecast results expected in the forthcoming PTA/SKA precision era [87, 88, 89, 90], which we will defer to future work. A Gaussian signal may turn out to be more significant in real data than is implied here, even though we suspect that the non-Gaussianity due to the noise will continue to dominate the cumulants. This is because the intrinsic noise parameters of each pulsar are measured, to an extent, post-sampling in the standard PTA search pipeline for a common spatially-correlated signal. The constrained noise parameters, particularly of the louder pulsars, may eventually be used to reduce their influence on the cumulants in both one- and two-point statistics. Real data is of course trickier, and there are plenty of challenges with PTA data that must be dealt with such as the practical limits to the sensitivity curve [58], and the fact that the timing data contains gaps.
On the other hand, the statistical properties of a Gaussian GWB in the one- and two-point function discussed in this work are general and independent of any particular primordial model. The non-Gaussianity in the one-point cumulants (left of Figure 2) and the two-point signal (3.1-3.4) are tied to the quadrupolar spatial correlation, due to the tensorial nature/polarization of GWs. This implies that analogous tests may be setup to study GWB in other GW bands, such as for future space based GW detectors that are expecting to meet a foreground GWB from white dwarfs, among other relevant cosmological GWBs motivated in theory. The formalisms drawn out in this work are easily extendible to accommodate non-Einsteinian and subluminal GW propagations [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102]. When the tests come to fruition, it will be inevitably important to relate constraints on non-Gaussianity in the GWB observables to theory, requiring predictions beyond the power spectrum [103, 104, 105]. We leave this to future work.
Data availability
Python notebooks that reproduce our results are published in GitHub.
Acknowledgments